OpenStudy (richyw):

Verify the identity in which \(\vec{a}\) is a constant vector, \(\vec{r}=x\hat{i}+y\hat{j}+z\hat{k}\), and \(r=||\vec{r}||\) \[\nabla r=\frac{\vec{r}}{r}\]

OpenStudy (richyw):

oops, damn brackets. (the bottom is what I need to verify)\[r=||\vec{r}||\]\[\nabla r=\frac{\vec{r}}{r}\]

OpenStudy (turingtest):

find each... what is \[\|\vec r\|\] ?

OpenStudy (richyw):

sorry I typed it wrong. Verify the identity in which a⃗ is a constant vector, \(\vec{r}=x\hat{i}+y\hat{j}+z\hat{k}\) and \(r=||\vec{r}||\) \[\nabla r=\frac{\vec{r}}{r}\] and ||r|| is simply the length of r

OpenStudy (turingtest):

I'm not sure I understand the question

OpenStudy (richyw):

Basically I need to verify that \[\nabla r=\frac{\vec{r}}{||r||}\]

OpenStudy (anonymous):

I believe some vector Geometry might help us here.

OpenStudy (anonymous):

 So if we divide the vector (and therefore also it's components) we obtain a unit vector of length one, just in an altered direction.